# Xing Huang a), Feng-Yu Wang a)

### v2

@@ 3. $(3)$
By (3.10) and (3.12), it suffices to prove that for any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$,
(3.58) $\begin{split}\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.\end{split}$
Let $R_{t}^{\varepsilon}$ be in (3.39). By (3.41) for $f$ replacing $V$, we obtain
(3.59) $\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}\big{[}f(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)\big{]},\ \ t\in(0,T].$
@@
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}=0,$
while $(B_{2})$, Lemma 3.5 and (3.39) lead to
$\lim_{\varepsilon\rightarrow 0}\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
in $L^{2}(\mathbb{P})$, by taking $\varepsilon\rightarrow 0$ in (3.59) we deduce (3.58) for $f\in C_{b}(\mathbb{R}^{d})$. By an approximation argument as in [29, Proof of (2.3)] for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, this implies (3.59) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$

### v4

@@ 2. $(2)$
By (3.12) and (3.13), it suffices to prove that for any $t\in(0,T]$ and $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$,
(3.52) $\begin{split}&\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
Let $R_{t}^{\varepsilon,x}$ be in (3.35). By (3.3) for $f$ replacing $U_{t}(y,\nu,\cdot)$, we obtain
(3.53) $\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}\mathbb{E}\big{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})(R_{t}^{\varepsilon,x}-1)\big{]}\mu(\text{\rm{d}}x),\ \ t\in(0,T].$
@@
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{x+\varepsilon\phi(x),\mu}-X_{t}^{x,\mu}|^{k}\Big{]}=0,$
while (3.43), (3.42), Lemma 3.5, (3.29), (3.26), (3.17) and (3.31) lead to
$\lim_{\varepsilon\rightarrow 0}\frac{R_{t}^{\varepsilon,x}-1}{\varepsilon}=\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}$
in $L^{2}(\mathbb{P})$, by taking $\varepsilon\rightarrow 0$ in (3.53) and using the dominated convergence theorem, we deduce (3.52) for $f\in C_{b}(\mathbb{R}^{d})$. By an approximation argument as in [30, Proof of (2.3)] for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, this implies (3.52) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$ By the approximation argument used in the proof of (3.4), we may further extend (3.52) to $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).$